OpenStudy (anonymous):

the slope of a line is -1/3. The point (-12, -6) is on the line. Determine the equation of this line

OpenStudy (anonymous):

\[y-y _{0}=m(x-x _{0})\] Where m is your slope, and (x0,y0) is your known point. This gives:\[y-(-6)=\frac{-1}{3}(x-(-12))\]Or,\[y+6=\frac{-1}{3}(x+6)\]Solving for y:\[y=\frac{-1}{3}x-2-6=\frac{-1}{3}x-8\]

OpenStudy (anonymous):

ooops screwed up the arithmitic. Should be y+6=(-1/3)x-4 which gives y=(-1/3)x-10

OpenStudy (anonymous):

In changed the 12 to a 6 for some reason :/